Kristin spent $131 on shirts. fancy shirts cost $28 and plain shirts cost $15. If she bought a total of seven how many of each kind did she buy?
Answer by tiffany222(56)

Let plain shirts be x
Let fancy shirts be y
x + y = 7
15x + 28y = 131
Multiply first equation by 28 to eliminate y:
28x + 28y = 196
15x + 28y = 131
Subtracting the two equations:
13x = 65
x = 5
Substituting x to get y:
x + y = 7
5 + y = 7
y = 2
Final Answer: 5 plain shirts and 2 fancy shirts